OpenStudy (anonymous):

symmetry of F(x)= x^2-4/x^2-2x

OpenStudy (anonymous):

f(x)=x²+2x-8 is in the form ax² + bx + x To find the x intercepts, factorise the equation and let y = 0: x² + 2x - 8 = (x + 4)(x - 2) = 0 x intercepts = -4 and 2 To find y intercepts let x = 0. x² + 2x - 8 0² + 2(0) - 8 = -8 y intercept = -8 The formula for axis of symmetry is -b/2a. b = 2, a = 1 -2/2 = -1 axis of symmetry = -1 Vertex occurs at the axis of symmetry. The axis of symmetry is the x coordinate of the vertex. To find the y coordinate, sub the value of the axis of symmetry into the equation. axis of symmetry = -1 y = x² + 2x - 8 y = -1² + 2(-1) - 8 y = 1 - 2 - 8 y = -9 turning point = (-1, -9) The domain is the values of x you can put into the equation. For this, every real number can be put in the equation, so the domain is R (all real numbers), expressed as: Domain = R The range are the y values that come out from putting any x values into the equation. The vertex is the lowest y value produced. The graph goes from -9 (the y coordinate of the turning point) upto infinity. The range is expressed as: Range = [-9, ∞)

OpenStudy (anonymous):

hope this helps ;o

OpenStudy (anonymous):

you didnt use my function tho...

OpenStudy (anonymous):

this was suppose to give you an example .. but hold on ill answer it for you (:

OpenStudy (anonymous):

ok thanks